# Manipulation Priors Learned From Egocentric Videos

### v1

@@ 6 Acknowledgements
Self-Contact. The first type of categorized error is self-contact, which occurs when the hand makes contact with the human body or worn clothing rather than an interacted object. In such cases, since the hand still appears to establish contact with a surface in the input image, the hand-object interaction segmenter [10] may misclassify these frames as contact frames. An example of a self-contact error is illustrated in Figure S13. A large proportion of training data exhibiting self-contact errors may induce an undesirable bias towards human garments in the encoder’s features. We hypothesize that this type of error can be tackled effectively by prompting visual grounding models [27, 29] for masks of people in the prediction frame and eliminating samples with contact points inside such masks.
@@
Figure S13: Self-Contact Error. The right hand is hovering above the actor’s shirt in (b), which is detected as a contact frame by the HOS segmenter.
@@
Figure S14: Premature Contact Error. The actor is reaching out for the garden scissors in (a), yet the hand is already detected as being in contact with the jug by the HOS segmenter, resulting in erroneous contact points in the prediction frame.
@@
S4.2 Error Analysis of Prediction Frame Search
@@
Figure S15: Premature Termination Error. In the prediction frame, the hand is still in contact with the stick.
@@
S4.3 Tokenizer Visualizations
Figure S16: Original and Tokenized Hand Meshes. The reconstructions (right of dashed lines) differ only slightly from the original input (left of dashed lines), preserving the original grasp.
We present examples of direct comparisons between original hand poses as regressed from Ego4D, and the hand poses after tokenization, in Figure S16. We find that our tokenizer learns to tokenize the hand poses well in most cases, with only small deviations observable between the original and the reconstructed data after tokenization.

### v2

@@ 6 Acknowledgements
Self-Contact. The first type of categorized error is self-contact, which occurs when the hand makes or appears to make contact with the human body or worn clothing rather than an interacted object. In such cases, since the hand still appears to establish contact with a surface in the input image, the hand-object interaction segmenter (HOS [EKVisor]) may misclassify these frames as contact frames. An example of a self-contact error is illustrated in Figure S18. A large proportion of training data exhibiting self-contact errors may induce an undesirable bias towards human garments in the encoder’s features. We hypothesize that this type of error can be tackled effectively by prompting visual grounding models [GLIP, GroundingDINO] for masks of people in the prediction frame and eliminating samples with contact points inside such masks.
@@
Figure S18: Visualization of Self-Contact Case. The right hand is hovering above the actor’s shirt in (b), which is detected as a contact frame by the HOS segmenter.
@@
Figure S19: Visualization of Premature Contact Case. The actor is reaching out for the garden scissors in (a), yet the hand is already detected as being in contact with the jug by the HOS segmenter, resulting in erroneous contact points in the prediction frame.
@@
S6.2 Error Analysis of Prediction Frame Search
@@
Figure S20: Visualization of Premature Termination Case. In the prediction frame, the hand is still in contact with the stick.
@@
S6.3 Tokenizer Visualizations
Figure S21: Visualization of Original and Tokenized Hand Meshes. The reconstructions differ only slightly from the original input, preserving the original grasp.
We present examples of direct comparisons between original hand poses as regressed from the Ego4D, and the hand poses after tokenization, in Figure S21. We find that our tokenizer learns to tokenize the hand poses well in most cases, with only small deviations observable between the original and the data reconstructed after tokenization.